# On Geometric Structures for Policy Parameterization in Continuous Control

### v2

@@ 5 Conclusion
This work demonstrates that effective continuous control does not require complex probability distributions. GAC achieves competitive or superior performance across diverse benchmarks using a single geometric operation: $\mathbf{a}=r\cdot\text{normalize}(w(\kappa)\cdot\bm{\mu}+(1-w(\kappa))\cdot\bm{\xi})$. By replacing distributional modeling with direct geometric mixing, we reduce parameter count by nearly 50% and improve performance, achieving a 37.6% improvement over SAC on Ant-v4 and the best results on 4 out of 6 MuJoCo benchmark tasks.
The success of GAC validates a broader principle: respecting the geometric structure of action spaces can be more effective than sophisticated probabilistic machinery. The consistent performance with a fixed radius $r=2.5$ across diverse tasks suggests that learning correct action directions is often the primary challenge, while magnitude can be addressed with simple task-specific constants.
@@
Limitations and Future Work. While GAC demonstrates strong empirical performance, several avenues remain for investigation. The action radius $r$, though robust across tasks (with $r=2.5$ working well for most benchmarks), could potentially be learned adaptively for task-specific optimization. The theoretical connection between our geometric exploration mechanism and information-theoretic quantities, while empirically validated through strong performance, warrants deeper mathematical analysis. Additionally, extending the geometric approach to discrete or hybrid action spaces presents exciting challenges for general-purpose control.
Despite these open questions, GAC’s success suggests that the Geometric Simplicity Principle—replacing probabilistic complexity with geometric structure—could transform other areas of RL. Future work might explore geometric approaches to value function approximation, hierarchical control, or multi-agent coordination. By showing that simple geometric operations can replace complex probabilistic frameworks, GAC challenges the prevailing assumption that sophisticated distributions are necessary for continuous control and opens new directions in geometric RL. Ultimately, control is not about predicting densities, but about choosing directions. GAC shows that when geometry is respected, simplicity is not a compromise—but a strength. We hope this work inspires further efforts to rethink RL through the lens of geometric structure, not just statistical modeling.
@@ A.1 Proof of Theorem 1
This result holds exactly for any dimension $d\geq 2$. ∎
@@ A.5 SAC Convergence with GAC
$L_{\text{actor}}=\mathbb{E}_{s\sim\mathcal{D},\bm{\xi}}[-\kappa(s)-Q(s,\mathbf{a})],$ (22)
where $\mathbf{a}$ is generated using GAC’s mechanism with random noise $\bm{\xi}$. This objective drives the policy toward high-value regions (via the Q-term) while maintaining exploration (via the $\kappa$ term), achieving similar goals to SAC’s entropy-regularized policy improvement without explicit distributional computations.
@@
Exploration–Exploitation Balance. The loss function induces an implicit tradeoff between exploration and exploitation. The direct gradient $\partial\mathcal{L}/\partial\kappa=-1$ promotes exploration by reducing $\kappa$, increasing action noise. In contrast, the Q-value term encourages exploitation: when higher $\kappa$ leads to better actions (and thus higher $Q$), gradients through $Q(s,\mathbf{a}(s;\kappa))$ push $\kappa$ upward. This dynamic balance emerges naturally from actor–critic interplay, without explicit entropy terms or temperature tuning.
@@ Appendix B Gradient Flow Analysis in Tanh-Squashed Policies
  • A substantial fraction (46.4%) of pre-squashed samples have gradients below 0.1, indicating severe saturation.

### v3

@@ 5 Conclusion
This work demonstrates that effective continuous control does not require complex probability distributions. GAC achieves competitive or superior performance across diverse benchmarks using geometric operations on the unit sphere: $\mathbf{a}=r\cdot\text{normalize}(w(\kappa)\cdot\bm{\mu}+(1-w(\kappa))\cdot\bm{\xi})$. By replacing distributional modeling with direct geometric mixing, we reduce parameter count by 50% while improving performance, achieving 37.6% gains over SAC on Ant-v4 and 112% on quadruped-run, with competitive or best results on 9 out of 12 tasks across MuJoCo and DMControl benchmarks.
The success of GAC validates a broader principle: respecting the geometric structure of action spaces can be more effective than sophisticated probabilistic machinery. Our fixed-radius design ($r=2.5$) works well across symmetric locomotion tasks by learning correct action directions, while the learnable per-dimension scaling variant (Eq. 6) provides additional flexibility for asymmetric scenarios, demonstrating GAC’s adaptability across diverse control challenges.
@@
Limitations and Future Work. While GAC demonstrates strong empirical performance, several avenues remain for investigation. The geometric prior (spherical manifold) proves particularly effective for coordinated locomotion but shows limitations on asymmetric tasks (e.g., Hopper’s single-leg dynamics, Pusher’s contact-rich manipulation). Future work could explore adaptive geometric structures that interpolate between spherical, elliptical, and unconstrained manifolds based on task characteristics. The theoretical connection between our geometric exploration mechanism and information-theoretic quantities, while empirically validated through strong performance, warrants deeper mathematical analysis. Additionally, extending the geometric approach to discrete or hybrid action spaces presents exciting challenges for general-purpose control.
Despite these open questions, GAC’s success suggests that the Geometric Simplicity Principle, which replaces probabilistic complexity with geometric structure, could transform other areas of RL.Future work might explore geometric approaches to value function approximation, hierarchical control, or multi-agent coordination. By showing that simple geometric operations can replace complex probabilistic frameworks, GAC challenges the prevailing assumption that sophisticated distributions are necessary for continuous control and opens new directions in geometric RL.
Ultimately, control is not about predicting densities, but about choosing directions. GAC shows that when geometry is respected, simplicity is not a compromise—but a strength. We hope this work inspires further efforts to rethink RL through the lens of geometric structure, not just statistical modeling.
@@ A.1 Proof of Theorem 1
This holds exactly for any $d\geq 2$, with no approximation error. ∎
@@ A.5 SAC Convergence with GAC
$L_{\text{actor}}=\mathbb{E}_{s\sim\mathcal{D},\bm{\xi}}[\kappa(s)-Q(s,\mathbf{a})],$ (23)
where $\mathbf{a}$ is generated using GAC’s mechanism with random noise $\bm{\xi}$. This objective drives the policy toward high-value regions (via the Q-Q term) while maintaining exploration (via the $\kappa$ term), achieving similar goals to SAC’s entropy-regularized policy improvement without explicit distributional computations.
@@
Exploration–Exploitation Balance. The loss function induces an implicit tradeoff between exploration and exploitation. The direct gradient $\partial\mathcal{L}/\partial\kappa=1$ promotes exploration by penalizing high $\kappa$, encouraging lower concentration and increased action noise. In contrast, the Q-value term encourages exploitation: when higher $\kappa$ leads to better actions (and thus higher $Q$), gradients through $Q(s,\mathbf{a}(s;\kappa))$ push $\kappa$ upward. This dynamic balance emerges naturally from actor–critic interplay, without explicit entropy terms or temperature tuning.
@@ Appendix B Gradient Flow Analysis in Tanh-Squashed Policies
  • A substantial fraction (46.4%) of pre-squashed samples have gradient magnitudes $|\tanh^{\prime}(\tilde{\mathbf{a}})|<0.05$, indicating severe saturation.
@@ C.3 Code and Model Release
Appendix D Adaptive Action Scaling Analysis
To validate the effectiveness of adaptive per-dimension scaling (Eq. 6), we analyze the learned action scales $\mathbf{r}\in\mathbb{R}^{d}$ from two complementary perspectives: (1) cross-task convergence patterns across the DMControl suite, and (2) ablation study on a canonical locomotion task.
D.1 Cross-Task Scale Convergence
Figure A.3: Training dynamics of learned action scales across DMControl tasks. Each line represents the mean scale $\bar{r}$ for one task (averaged over all action dimensions and 5 seeds). Scales converge to task-specific values in $[1.0,3.0]$ after initial exploration (<<100K steps), demonstrating adaptive learning without manual tuning.
Figure A.3 shows the evolution of mean scales $\bar{r}=\frac{1}{d}\sum_{i=1}^{d}r_{i}$ across 6 DMControl tasks over 500K steps.
Key Observations:
1) Stable Convergence Across Tasks. After an initial exploration phase (<<100K steps), all tasks exhibit stable scale convergence, indicating the geo-head successfully identifies appropriate action magnitudes without manual intervention.
2) Task-Dependent Adaptation. Different morphologies learn distinct scale profiles:
  • Walker tasks (walk, run): Converge to $\bar{r}\approx 2.0$$2.3$, close to our fixed $r=2.5$ baseline, confirming the validity of uniform scaling for symmetric bipedal locomotion.
  • Quadruped tasks: Stabilize at $\bar{r}\approx 2.0$$2.5$, with more variance reflecting multi-leg coordination requirements.
  • Cheetah-run: Tight convergence to $\bar{r}\approx 2.3$, consistent with forward-acceleration-dominant dynamics.
  • Fish-upright: Gradual rise from $\approx 1.0$ to $\approx 2.2$, balancing buoyancy control and orientation stabilization.
3) Geometric Validity. The learned range $[1.0,3.0]$ aligns with geometric constraints: since normalized directions satisfy $|\hat{v}_{i}|\leq 1$, scales $r_{i}\in[1.0,3.0]$ produce per-dimension actions $|a_{i}|\lesssim 3.0$, comfortably within normalized bounds $1,1]\times\text{action\_limit}$ for MuJoCo/DMControl, avoiding both under-actuation ($r<1.0$) and over-saturation ($r>3.0$).
D.2 Ablation: Fixed vs. Adaptive Scaling on Walker2d-v4
To rigorously validate the necessity of learnable $\mathbf{r}$ versus fixed $r$, we conduct a controlled ablation on Walker2d-v4, comparing:
  • Fixed $r\in\{1.0,2.0,3.0\}$ (scalar radius)
  • Adaptive $\mathbf{r}\in\mathbb{R}^{6}$ (learned per-dimension)
All experiments use 5 random seeds with identical hyperparameters except for the scaling strategy.
(a) Learning curves
(b) Adaptive $\bar{r}$ evolution
(c) Adaptive $\text{std}(r_{i})$ evolution
Figure A.4: Ablation on Walker2d-v4: Fixed vs. Adaptive Scaling. (a) Learning curves show adaptive $\mathbf{r}$ achieves the best final performance (5064 $\pm$ 156) but only marginally outperforms fixed $r=2.0$ (4607) and $r=3.0$ (4919), while $r=1.0$ (3602) significantly underperforms due to insufficient action magnitude. (b) The learned mean scale $\bar{r}$ converges to $\approx 1.7$, between the $r=1.0$ and $r=2.0$ baselines. (c) The standard deviation $\text{std}(r_{i})$ rises from $\approx 0.7$ to $\approx 0.9$, indicating dimension-specific differentiation rather than uniform scaling collapse.
Results & Implications:
1) Marginal Gain on Symmetric Tasks. Table A.2 shows adaptive $\mathbf{r}$ achieves only 3–9% improvement over well-tuned fixed $r\in\{2.0,3.0\}$, confirming our hypothesis that symmetric locomotion does not require per-dimension scaling. The small gain comes from fine-tuning rather than fundamentally different geometry.
Table A.2: Final performance (500K steps) on Walker2d-v4
Method Return Params
$r=1.0$ (fixed) $3602\pm 130$ 0
$r=2.0$ (fixed) $4607\pm 179$ 0
$r=3.0$ (fixed) $4920\pm 250$ 0
Adaptive $\mathbf{r}$ $\mathbf{5064\pm 156}$ +384+384 (geo-head)
2) Low Sensitivity of Fixed $r$. The $r=2.0$ and $r=3.0$ baselines differ by less than 7%, demonstrating that GAC is not sensitive to radius hyperparameters within a reasonable range. This validates the robustness of our default choice $r=2.5$ in the main paper (Section 3.5).
3) Learned Scales Confirm Theoretical Design. Figure 4(b) shows adaptive $\bar{r}$ converges to $\approx 1.7$, consistent with the $[1.0,3.0]$ range observed in DMControl (Figure A.3). The increasing $\text{std}(r_{i})$ (Figure 4(c)) indicates the network learns dimension-specific modulation rather than collapsing to uniform scaling, though the impact is modest for this symmetric task.
4) Task Complexity Determines Necessity. Combining this ablation with DMControl results (Section 4.2), we conclude:
  • Simple/symmetric tasks (Walker2d, HalfCheetah): Fixed $r$ suffices, saving 50% parameters and avoiding unnecessary complexity.
  • Complex/asymmetric tasks (Quadruped, high-DoF manipulation): Adaptive $\mathbf{r}$ unlocks 6–112% gains by enabling fine-grained per-dimension control.
This demonstrates that $r$ is not a hyperparameter requiring tedious tuning, but rather an environment-dependent geometric parameter that can be either fixed (for efficiency) or learned (for flexibility) based on task structure.
D.3 Summary
This two-level analysis validates GAC’s design philosophy:
  • Macro-level (DMControl): Adaptive $\mathbf{r}$ converges robustly across diverse morphologies, with task-specific profiles emerging naturally without manual tuning.
  • Micro-level (Walker2d): On symmetric tasks, fixed $r$ achieves near-optimal performance with fewer parameters, while adaptive $\mathbf{r}$ provides modest gains at the cost of added complexity.
Together, these findings demonstrate GAC’s flexibility: structural simplicity when sufficient, adaptive complexity when necessary. The fixed $r=2.5$ baseline in the main paper represents an efficiency-optimized choice validated by both theoretical geometry and empirical convergence, while the GAC-Scale variant (Eq. 6) extends this framework to handle asymmetric scenarios without sacrificing stability.
Appendix E Computational Complexity Analysis
We provide a detailed comparison of per-sample computational costs across GAC, SAC, and TD3:
Operation Breakdown:
Table A.3: Computational complexity per action sample. GAC achieves efficiency comparable to deterministic methods while maintaining structured exploration.
Method Core Operations Cost per Sample
TD3 Forward pass + Gaussian noise $O(d)$
SAC Sampling + $\tanh$ + log-det Jacobian $O(d)$ with log overhead
GAC Normalization + spherical interpolation $O(d)$
Detailed Cost Analysis:
[…]
GAC: Requires (1) Direction generation (forward pass + L2 norm) [$O(d)$], (2) Spherical noise sampling (normalized Gaussian) [$O(d)$], and (3) Linear interpolation [$O(d)$]. Total: $3\times O(d)$ with lightweight operations (no logarithms or special functions).
TD3: Baseline deterministic policy [$O(d)$] plus Gaussian noise [$O(d)$]. Total: $2\times O(d)$.
Empirical Timing (NVIDIA RTX 3090):
  • SAC: 1.0× (baseline)
  • GAC: 0.78× (22% faster than SAC)
  • TD3: 0.65× (35% faster than SAC)
GAC achieves near-TD3 efficiency while maintaining the structured exploration benefits of stochastic policies, eliminating the computational overhead of entropy regularization that SAC requires.
Parameter Efficiency:
  • SAC: $2d$ outputs (mean $\mu$, log-std $\log\sigma$)
  • GAC (fixed $r$): $d+1$ outputs (direction $\bm{\mu}$, concentration $\kappa$) $\approx$ 50% reduction
  • GAC-Scale: $2d+1$ outputs (direction $\bm{\mu}$, concentration $\kappa$, scales $\mathbf{r}$) — comparable to SAC but with geometric structure
This analysis demonstrates that GAC’s geometric approach achieves computational efficiency comparable to deterministic methods while preserving the exploration benefits of stochastic policies.